solve by the substitution method.
9x+5y=-29
x=24-6y

Answer by mananth(16949)

9x+5y=-29 .............1
6y=24-x
/6 =
y=4-0.17 x
substitute y in (1)
9x+5(4-0.17x)=-29
9x+20-0.83x=-29
9x-0.83x =-29-20
8.17x=-49
/8.17
x=-6
Plug the value of x in (1)
9x+5y=-29
9*-6+5y= -29
-54+5y=-29
5y=25
/5
y=5

CHECK
6y=24-x
30=24+6